a grindstone of radius 6.1m is initially spinning with an angular speed of 3.4 radians per second. The angular speed is the increased to 8.9 radians per second over the next 4 seconds. Assume that the angular acceleration is constant. What is the average angular speed of the grindstone
1
Expert's answer
2019-11-11T15:51:40-0500

If the angular acceleration is constant the average angular speed of the grindstone is


"\omega_{av}=\frac{\omega_i+\omega_2}{2}"

"\omega_{av}=\frac{3.4+8.9}{2}=6.15\frac{rad}{s}"